Find the solution of this system of edquations

6x = 90 + 3y
X - 3y = 50

Answer by ewatrrr(10682)


Hi
6x = 90 + 3y
x - 3y = 50
6x -3y = -90
x - 3y = 50

6x - 3y = -90
-x + 3y = -50
5x = -140
x = -28 and y = (x-50)/3 = -78/3 = -26
Ordered pair (-28,-26) is the solution for this system
CHECKING our Answer***
6*-28 - 3*-26 = -168 + 78 = -90